# Evidence for Low Universal Equilibrium Black Hole Spin in Luminous Magnetically Arrested Disks

### v1

@@ 3.3.2 Decoupling of thermal and magnetic dynamics
We conclude that when the disk thickness falls below a certain threshold, the disk properties are determined more strongly by the disk magnetic field structure than by the disk thermodynamics. This is evidenced by the minimal differences in $F_{\rm b,{jet}}$, $\eta_{\rm EM}$, $k$, and $\frac{1}{2}\langle\delta b^{2}\rangle/(p_{\rm g}+p_{\rm rad})$ between simulations with $h/r=0.1$ and $h/r=0.03$. Consequently, we suggest that this is why the equilibrium spin, $a_{\rm eq}$, converges as the disk thickness decreases. For disk thickness values below $h/r\sim 0.1$, jet properties become decoupled from the disk thermodynamics, because magnetic forces dominate over thermal forces.
@@
Figure 11: All thin MADs, $0.03\leq h/r\leq 0.1$, experience comparable magnitude of magnetic compression. In contrast, for thick MADs ($h/r=0.3$), compression is twice as strong: $F_{\rm thick}\sim 2F_{\rm thin}$. This is consistent with thicker disks producing more powerful jets that compress the disk more. The density-normalized compression force, $F_{\rm b,{jet}}$, due to the large-scale magnetic field is shown as a function of $\theta$, defined in Eq. (27) and measured at $r=3\,\,r_{g}$. Red arrows indicate the force direction, with positive forces pointing along $\theta$ and negative forces pointing against it. We show the high resolution simulation for $h/r=0.3$ (H3a0.94hr).
3.4 Semi-analytic model for BH spin-down in cooled MADs
Lowell et al. (2024) developed a physically-motivated semi-analytic model that describes BH spin-down in radiatively-inefficient MADs. Using our new suite of simulations of luminous MADs with varying $h/r$, we now construct a model for thinner MADs.
In panels (a) of both Figures 4 and 5, we find that $e_{\rm HD}$ follows a curve of $0.85e_{\rm NT}$, and $l_{\rm HD}$ follows $0.4l_{\rm NT}$. Thus, we can model the hydrodynamic disk component of cooled MADs as follows: $s_{\rm HD,cool}=0.4l_{\rm NT}-0.85e_{\rm NT}\times 2a$.
The electromagnetic jet component also follows an interesting trend that is straightfoward to model. As we showed in Section 3.3, $\eta_{\rm EM}$ for $h/r=0.1$ follows a curve that is approximately $1/2$ of $\eta_{\rm EM}$ for thick MADs. Thus, we can use the fit from Lowell et al. (2024) and model $\rm\eta_{\rm EM}$ for thin disk as $\eta_{\rm EM}^{\rm thin}=\frac{1}{2}\eta_{\rm EM}^{\rm thick}$, where
$\eta_{\rm EM}^{\rm thick}\times 100=\left\{\begin{array}[]{ll}-19.8a^{4}+48.9a^{2},&a\leq 0,\\ 106.3a^{4}+39.5a^{2},&a\geq 0.\\ \end{array}\right.$ (28)
Figure 6 (top) shows that our model for $\rm\eta_{\rm EM}$ above well reproduces the jet efficiency for thin disks.
We fit the spin-dependence of the $k$ factor using the data shown in Fig. 7, with a power law  [^5] In this definition, we set $k(a=0)=0.2$, though strictly speaking, $k(a=0)$ is undefined. This choice, made for the quality of the fit, does not affect our results since the angular momentum extraction term goes to zero when $a=0$, as $\eta_{\rm EM}(a=0)=0$.,
$k_{\rm thin}=\left\{\begin{array}[]{ll}0.16\,|a|^{0.15}+0.2,&a<0\\ 0.24\,a^{0.4}+0.2,&a>0.\\ \end{array}\right.$ (29)
Figure 4, shows that our fit for $l_{\rm EM}=\frac{1}{k_{\rm thin}\Omega_{H}}\eta_{\rm EM}^{\rm thin}$, as a function of $a$, reproduces the simulation data well. By combining EM and HD components, we obtain the full semi-analytic model,
$\begin{split}s_{\rm MAD,thin}&=0.4\times l_{\rm NT}-0.85\times e_{\rm NT}\times 2a\\ &-\frac{\eta_{\rm EM}^{\rm thick}}{2}\left(\frac{1}{k_{\rm thin}\Omega_{\rm H}}-2a\right),\end{split}$ (30)
for BH spin-down in cool MADs. Figure 2 shows our spin evolution model (Eq. 30) with the purple dashed line, which reproduces the simulation data well. Although this model is tailored to $h/r=0.1$, it also well reproduces the behavior of $s$ for thinner disks, thanks to the universality of thin MAD structure (see Section 3.3).
3.5 BH spin-down timescales
Figure 12 shows BH spin evolution as a function of mass accreted onto the BH for thin MADs with $h/r=0.1$. For this, we solve the coupled ODEs, Equations (2) and (4), and use spline interpolations for the specific energy and angular momentum fluxes at the horizon, for simulations (solid lines) with $h/r=0.1$ (models H1a#). We show solutions for the initial spin values $a_{0}=-1$, 00, $0.5$, and $1$. We find that regardless of the initial spin, $a$ evolves to an equilibrium value of $a_{\rm eq}\approx 0.3$. Most initial spins equilibrate after the BH has accreted 85% of its initial mass, while $a_{0}=-1$ requires 125% accretion to reach equilibrium. Thin MADs with $a=1$ spin down to $a=0.5$ after accreting 25% of the initial BH mass, a slower process compared to thick MADs, which only require the accretion of 10% of the BH mass for the same spin-down. We also solve the ODEs using our thin MAD model (Equation 30, dashed lines) and confirm that our model reproduces the simulation (solid lines) well. We associate the reduced BH spin-down efficiency in thin MADs with a general decrease in angular momentum and energy extraction by the jets (see Section 3.3).
Figure 12: Our semi-analytic model (dashed lines) accurately describes BH spin evolution in thin MADs. BH spin evolution $a(m/M_{0})$ for thin MADs with $h/r=0.1$ (models H1a#). Here, $m$ is the accreted mass and $M_{0}$ is the initial mass of the BH. When $m/M_{0}=1$, the BH has accreted its own (initial) mass. We show solutions for the initial spin values $a_{0}=-1$, 00, $0.5$, and $1$, for any initial spin, the BH will evolve to an equilibrium spin of $a_{\rm eq}\approx 0.29$. For an initial spin of $a=1$, the BH of a thin MAD will spin down to $a=0.5$ after accreting $25\%$ of its initial mass. For most initial spins, the BH must accrete around $85\%$ to reach equilibrium spin. For negative initial spin, the BH must accrete more mass than the positive initial spin to reach $a_{\rm eq}$, roughly $125\%$ of $M_{0}$. Dashed lines, which show the spin evolution using the semi-analytic model (Eq. 30), agree well with thin the MAD spin-down model, which uses spline fits to $e_{\rm in}$ and $l_{\rm in}$.
4 Summary and Discussion
4.1 Summary
@@
In this study, we performed 3D GRMHD simulations of thin, luminous MADs with various disk thicknesses to investigate BH spin evolution. Previous works found that thick, nonradiative MADs result in low equilibrium spin values, $a_{\rm eq}\simeq 0.07$ for polytropic index $\Gamma=4/3$ (Tchekhovskoy et al., 2011; Lowell et al., 2024) and $a_{\rm eq}\simeq 0.035$ for $\Gamma=13/9$ (Narayan et al., 2022; Lowell et al., 2024). Our findings reveal, for the first time, that even in cooled MADs, the BH spin reduces down to a small equilibrium spin value of $a_{\rm eq}^{\rm thin}\simeq 0.3$. Interestingly, we observe that from $h/r=0.1$ to $h/r=0.03$, the difference in the equilibrium spin is negligible. This suggests that the equilibrium spin converges to a universal value, $a_{\rm eq}\approx 0.3$. This is in stark contrast to the equilibrium spin of $a_{\rm eq}\approx 1$ for the analytic NT disk. Ultimately, it is straightforward to achieve a low BH spin in a system that accretes a significant amount of mass in the MAD state, be it of the luminous or nonradiative variety. Here, we used the physical model of Lowell et al. (2024) to decompose the torques of cooled MADs and analyze their BH spin evolution. As with thick MADs, the jet electromagnetic torques remain the dominant spin-down mechanism. However, we find that the higher equilibrium spin, by a factor of four, in thin MADs ($h/r\leq 0.1$) relative to thick MADs ($h/r=0.3$) is due to less efficient jet-driven energy and angular momentum extraction from the BH. Specifically, jets in thin MADs are half as energetically efficient as those in thick MADs. Further, thin MADs have EM angular momentum torques up to two times weaker than thick MADs, for the same jet efficiency. We attribute this reduced efficiency in jet-driven angular momentum extraction to differences in jet shape between thin and thick MADs. Whereas thick MADs produce parabolic jets, thin MADs generate monopolar jets, which are less efficient at angular momentum extraction for the same jet energy efficiency.
@@
We present a new semi-analytic spin-down model that accurately reproduces the spin-down timescale and equilibrium spin of thin MADs with $h/r\leq 0.1$. We find that in a thin MAD, a BH with an initial spin of $a=1$ spins down to $a=0.5$ after accreting just $25\%$ of its initial mass. This is a factor of $2.5$ larger than for thick ($h/r=0.3$) MADs. Most initial spins reach equilibrium after accreting $85\%$ of the BH’s initial mass, while $a=-1$ equilibrates after accreting $125\%$.
@@
Our results indicate that some persistent MAD physical property ensures the universality of the equilibrium spin for thermal scale heights less than $h/r=0.1$. We propose that the explanation lies in the balance between thermal, radiation, and magnetic (large-scale or turbulent) pressures. Whereas thick MADs are thermally-supported, thin MADs (thermal $h/r\leq 0.1$) are supported by the turbulent magnetic pressure (as also found by Scepi et al. (2023)). Furthermore, we find that in such thin MADs the vertical compression of the disk by the large-scale magnetic field pressure, associated with the jet-launching, becomes independent of the disk thermal thickness. This suggests that the overall magnetic structure, which governs most of the BH spin-down, decouples from the accretion disk thermal properties for $h/r\leq 0.1$. In such thin MADs, the decoupling of thermal and magnetic pressures (large-scale or turbulent) may explain why the equilibrium BH spin settles at a universal, low value, $a_{\rm eq}\approx 0.3$, independent of the disk thickness.
@@
Our resolution tests ( Appendix A) show that under-resolving thin MADs with $h/r=0.05$ leads to a significantly higher equilibrium spin. We find that a vertical resolution of $\tilde{N}_{\theta}\lesssim 5$, where $\tilde{N}_{\theta}$ is the number of cells per scale height, is insufficient to accurately determine the equilibrium spin  [^6] Difference in $\varphi$-resolution between low- and high- resolution simulation might also affect the equilibrium spin.. This discrepancy arises from the artificially increased hydrodynamic energy and angular momentum fluxes in lower-resolution runs, leading to $a_{\rm eq}\approx 0.4$ for $\tilde{N}_{\theta}\approx 5$ and $a_{\rm eq}\approx 0.3$ for $\tilde{N}_{\theta}\approx 9$. We emphasize that high resolution is essential for accurately modeling spin evolution in thin MADs.
@@
It is possible that for even thinner disks, perhaps an order of magnitude thinner in $h/r$, magnetic and thermal pressures remain permanently decoupled. Alternatively, disk dynamics may still depend on $h/r$, reaching a critical value where the disk can no longer retain its magnetic flux (Lubow & Spruit, 1995). In that scenario, the disk would transition toward the NT state, theoretically contributing to BH spin-up. Observational evidence, such as jet quenching in the razor-thin XRB soft state, may support this idea. In contrasst to thin MADs, which are dominated by super-thermal mainly turbulent magnetic fields at the midplane accretion disks dominated by super-thermal toroidal fields (which is different from MADs) appear to be influenced by their thermal thickness, even if $p_{\rm g}$ is not dynamically relevant for the disk vertical equilibrium (Squire et al., 2024).
@@
We compute a simple fit for the dependence of $a_{\rm eq}$ on disk thickness, finding $a_{\rm eq}\simeq 0.31-2.7(h/r)^{2}.$ This provides a rough estimate of how $a_{\rm eq}$ might vary with the Eddington ratio, $\lambda_{\rm Edd}=\dot{m}/\dot{m}_{\rm Edd}$, and clearly shows that the equilibrium spin plateaus at $a_{\rm eq}\lesssim 0.31$. In future work, we plan to extend this model across a broader range of Eddington ratios.
4.2 Comparison with other work
@@
Although our work is the first to measure BH spin-down for cooled thin MADs with $h/r\leq 0.1$, previous studies have already examined jet efficiency in such disks (Avara et al., 2016; Scepi et al., 2023; Dhang et al., 2024). For $h/r=0.1$, our results broadly agree with prior findings, showing similar jet efficiencies and magnetic fluxes at the event horizon. However, for even thinner disks ($h/r\sim 0.03$), our measured jet energy efficiencies differ significantly, with Scepi et al. (2023) reporting jet efficiencies a factor of 5 lower. This discrepancy may stem from the differences in the measurement methods, as Scepi et al. (2023) evaluates jet power at larger distances rather than the event horizon. Nonetheless, we note that our measurement of magnetic flux at the horizon, $\phi_{\rm BH}$, is consistent with Scepi et al. (2023), even for MADs with $h/r\sim 0.03$.
@@
Recent work by Ricarte et al. (2023) measured equilibrium spin in radiative MAD simulations but focused on thicker disks, with a minimum [^7] We note that our measurements of $h/r$ are not entirely comparable, as they use radiative and thermal pressure to measure $h/r$. $h/r\sim 0.18$. Direct comparison is difficult since our radiative simulations do not overlap in Eddington ratio, $\lambda_{\rm Edd}$, space. However, for $h/r\sim 0.18$, they report a higher equilibrium spin ($a_{\rm eq}\sim 0.8$) than what we find for $h/r=0.1$ ($a_{\rm eq}\sim 0.3$). Within our model, we expect $0.07\leq a_{\rm eq}\leq 0.3$ for such disks, in tension with their results.
@@
Ricarte et al. (2023) also propose a general spin-down model applicable to all Eddington ratios, which we can compare to our radiative simulations at $\lambda_{\rm Edd}=0.35$ which have $h/r=0.03$ (Liska et al., 2022). At this Eddington ratio they report an equilibrium spin of $a_{\rm eq}\sim 1$, whereas we find $a_{\rm eq}\sim 0.3$. Moreover, their results show no sign of convergence in $a_{\rm eq}$ for thinner disks; instead, their equilibrium spin continues increasing as disk thickness decreases, eventually reaching values consistent with the NT disk ($a_{\rm eq}=1$).
4.3 Observational implications
@@
This work is timely for understanding BH spin measurements in binaries, where mass transfer—the process by which a donor star loses mass to a BH—is a key aspect of evolution and may also govern the BH spin evolution. Stellar-mass BH spins are directly measured via gravitational waves (GW) in binary mergers (Abbott et al., 2019; Wysocki et al., 2019; LIGO Scientific Collaboration et al., 2023; Edelman et al., 2023) and constrained in XRBs using X-ray reflection (Liu et al., 2008; Draghis et al., 2024). However, these methods yield conflicting spin distributions: GW observations show a distribution centered around $|a|\sim 0.2$, whereas EM observations find a distribution skewed toward higher BH spins, $a\gtrsim 0.5$. This discrepancy may arise from GW sources and XRBs belonging to distinct binary populations, which have different mass transfer histories (Gallegos-Garcia et al., 2022; Fishbach & Kalogera, 2022). Spin evolution during mass transfer involves violent accretion episodes that can significantly alter the BH natal spin (Qin et al., 2022). Whether these hypercritical accretion phases are linked to ultraluminous X-ray sources (ULXs) remains uncertain, but the mere fact that ULXs exist, confirms that some binaries accrete far above Eddington (King et al., 2023). However, binary evolution studies typically assume the NT accretion disk model for spin evolution, which does not apply to disks with large-scale vertical magnetic fields. Our MAD spin-down model could yield significantly different spin distributions compared to NT models, making it a valuable case study for spin evolution in binary evolution codes.
@@
A leading model for long GRBs is the collapsar scenario, where a rapidly rotating massive star collapses into a BH, launching powerful jets that produce gamma-ray emission (Woosley, 1993). Our previous work on thick MADs (Lowell et al., 2024) applied to collapsar GRBs demonstrated that their high accretion rates and strong jets make them ideal systems for BH spin-down (Jacquemin-Ide et al., 2024a; Wu et al., 2024). We showed that collapsars can spin down to $a_{\rm eq}=0.07$ within seconds, well before the GRB ends. However, this work did not account for neutrino cooling (Chevalier, 1989), which alters disk thickness and, consequently as shown here equilibrium spin. Indeed, recent work by Issa et al. (2025) finds that under neutrino cooling, collapsars spin down to a higher equilibrium spin, $a_{\rm eq}\simeq 0.13$, and therefore fall between our cases of $h/r=0.1$ and $h/r=0.3$ consistent with our results. Their equilibrium spin measurement could serve as an initial condition for BH spin, which our model—better suited for binary evolution—would then further evolve.
@@
Jet feedback is crucial for galaxy formation and the evolution of large-scale cosmological structures like galaxy clusters. Understanding BH spin evolution in luminous quasars is essential, as BH spin changes directly impact the energy injection into the galaxy and cluster environement. Recent quasar formation simulations, which track accretion disk formation and evolution within a dynamically simulated galaxy, suggest the emergence of a superthermal, toroidally dominated disk (Hopkins et al., 2024). By interpolating the final snapshot of Hopkins et al. (2024) simulation to H-AMR, Kaaz et al. (2025) found that the central BH can accrete at trans-Eddington rates, $\lambda_{\rm Edd}\equiv\dot{m}/\dot{m}_{\rm Edd}\sim 5$. The disk settles into a MAD-like structure similar to the thin MADs studied here. Their simplified spin-down estimates suggest spin evolution on timescales comparable to the AGN duty cycle, $\sim 10\,\,\rm{Myr}$. This indicates that our results on BH spin-down are highly relevant for cosmological quasar accretion, as these systems accrete rapidly and exhibit MAD-like structures. This is especially important if one wants to also understand the implications of AGN BH spin measurements for the BH accretion and merger history (Reynolds, 2021).
@@
Observations of jetted accretion disk sources (XRBs and AGN) reveal a strong correlation between jet luminosity ($L_{\rm jet}$) and accretion luminosity ($\dot{m}$), independent of accretor mass, and spanning a wide range of Eddington ratios ($\lambda_{\rm Edd}$) (Serjeant et al., 1998; Corbel et al., 2003; Markoff et al., 2003; Corbel et al., 2013; Zamaninasab et al., 2014). This suggests little variation in jet efficiency, $\eta_{\rm EM}=L_{\rm jet}/\dot{m}c^{2}$, across these systems. Our results show that in MADs, $\eta_{\rm EM}$ remains independent of disk thickness for $h/r\leq 0.1$, and even when comparing thick ($h/r=0.3$) and thin ($h/r=0.1$) disks, jet efficiency decreases by only a factor of 2 for fixed $a$. The small variation, within a factor of 2, in $\eta_{\rm EM}$ for fixed BH spin across different disk properties would be hard to distinguish observationally. This might explain the universal correlation between jet luminosity and accretion rate observed across diverse astrophysical systems and Eddington.
@@
Our results suggest that if all jetted BHs are in the MAD state, they should spin down to low spin values, $a_{\rm eq}\lesssim 0.3$, regardless of their Eddington ratio, given sufficient time. This has significant implications for the cosmological evolution of BHs in AGN, directly impacting cosmological BH feedback. Additionally, in binary systems, these findings are crucial for LISA and LVK science, offering a potential explanation for the observed BH spin distribution.

### v2

@@ V.2 Decoupling of thermal and magnetic forces
We conclude that when the disk thickness falls below a certain threshold, the disk properties are determined more strongly by the disk magnetic field structure than by the disk thermodynamics. This is evidenced by the minimal differences in $F_{\mathrm{b,{jet}}}$, $\eta_{\mathrm{EM}}$, $k$, and $\frac{1}{2}\langle\delta b^{2}\rangle/(p_{\mathrm{g}}+p_{\mathrm{rad}})$ between the simulations with $h/r=0.1$ and $h/r=0.03$. Consequently, we suggest that this is why the equilibrium spin, $a_{\mathrm{eq}}$, converges as the disk thickness decreases. For disk thickness values below $h/r\sim 0.1$, jet properties become decoupled from the disk thermodynamics, because magnetic forces dominate over thermal forces.
@@
Figure 12: All thin MADs, $0.03\leq h/r\leq 0.1$, experience comparable magnitude of the magnetic compression force, $F_{b,\rm jet}$, due to the twin polar jets, as shown here for BH spin, $a=0.94$. In contrast, for thick MADs ($h/r=0.3$), the compression force is twice as strong: $F_{\mathrm{thick}}\sim 2F_{\mathrm{thin}}$. This is consistent with thicker disks producing more powerful jets that compress the disk more. The density-normalized compression force, $F_{\mathrm{b,{jet}}}$, due to the large-scale magnetic field is shown as a function of $\theta$, defined in Eq. (36) and measured at $r=3\,\,r_{\mathrm{g}}$. Red arrows indicate the force direction, with positive forces pointing along $\theta$ and negative forces pointing against it. We show the high-resolution simulation for $h/r=0.3$ (H3a0.94hr).
VI Summary and Discussion
VI.1 Summary
@@
In this study, we performed 3D GRMHD simulations of thin, luminous MADs with various disk thicknesses to investigate BH spin evolution in time. Previous works found that thick nonradiative MADs ($h/r\simeq 0.3$) result in low equilibrium spin values, $a_{\mathrm{eq}}\simeq 0.07$ for gas polytropic index $\Gamma=4/3$ [92, 96, 54] and $a_{\mathrm{eq}}\simeq 0.035$ for $\Gamma=13/9$ [65, 54]. Our findings reveal, for the first time, that even in thin MADs ($h/r\leq 0.1$) the BH spins down to a small equilibrium spin value, $a_{\mathrm{eq}}^{\mathrm{thin}}\simeq 0.3$. Interestingly, we observe that from $h/r=0.1$ to $h/r=0.03$, the difference in the equilibrium spin is negligible. This suggests that the equilibrium spin converges to a universal value $a_{\mathrm{eq}}\approx 0.3$. This is in stark contrast to the equilibrium spin of $a_{\mathrm{eq}}\approx 1$ for the analytic NT disk. Ultimately, it is straightforward to achieve a low BH spin in a system that accretes a significant amount of mass in the MAD state, be it of the luminous or nonradiative variety. Here, we used the physical model of Lowell et al. [54] to decompose the torques of thin MADs and analyze their BH spin evolution. As with thick MADs, the jet electromagnetic torques remain the dominant spin-down mechanism. However, we find that the higher equilibrium spin, by a factor of $4$, in thin MADs ($h/r\leq 0.1$) relative to thick MADs ($h/r=0.3$) is due to less efficient jet-driven energy and angular momentum extraction from the BH. Specifically, jets in thin MADs are half as energetically efficient as those in thick MADs. Further, thin MADs have EM angular momentum torques up to $2$ times weaker than thick MADs, for the same jet efficiency. We attribute this reduced efficiency in jet-driven angular momentum extraction to differences in jet shape between thin and thick MADs. Whereas thick MADs produce more parabolic jets, thin MADs generate comparatively more monopolar jets, which are less efficient at angular momentum extraction for the same jet energy efficiency.
@@
We present a new semi-analytic spin-down model that accurately reproduces the spin-down timescale and equilibrium spin of thin MADs with $h/r\leq 0.1$. We find that in a thin MAD, a BH with an initial spin of $a=1$ spins down to $a=0.5$ after accreting just $25\%$ of its initial mass. This is a factor of $2.5$ larger than for thick ($h/r=0.3$) MADs. Most initial spins reach equilibrium after accreting $85\%$ of the BH’s initial mass, while a=1a=-1 equilibrates after accreting $125\%$.
@@
Our results indicate that some persistent MAD physical property ensures the universality of the equilibrium spin for thermal scale heights less than $h/r=0.1$. We propose that the explanation lies in the balance between thermal, radiation, and magnetic (large-scale or turbulent) pressures. Whereas thick MADs are thermally supported, thin MADs (of thermal $h/r\leq 0.1$) are supported by the turbulent magnetic pressure (as also found by Scepi et al. [84]). Furthermore, we find that in such thin MADs the vertical compression of the disk by the large-scale magnetic field pressure, associated with the jet launching, becomes independent of the disk thermal thickness. This suggests that the overall magnetic structure, which governs most of the BH spin-down, decouples from the accretion disk thermal properties for $h/r\leq 0.1$. In such thin MADs, the decoupling of thermal and magnetic pressures (large-scale or turbulent) may explain why the equilibrium BH spin settles at a universal, low value, $a_{\mathrm{eq}}\approx 0.3$, independent of the disk thickness.
@@
Our resolution tests (Appendix A) show that under-resolving thin MADs with $h/r=0.05$ leads to a significantly higher equilibrium spin. We find that a vertical resolution of $\tilde{N}_{\theta}\lesssim 5$, where $\tilde{N}_{\theta}$ is the number of cells per scale height, is insufficient to accurately determine the equilibrium spin  [^7] Difference in $\varphi$-resolution between low- and high-resolution simulation might also affect the equilibrium spin.. This discrepancy arises from the artificially increased hydrodynamic energy and angular momentum fluxes in lower resolution runs, leading to $a_{\mathrm{eq}}\approx 0.4$ for $\tilde{N}_{\theta}\approx 5$ and $a_{\mathrm{eq}}\approx 0.3$ for $\tilde{N}_{\theta}\approx 9$. We emphasize that high resolution is essential for accurately modeling spin evolution in thin MADs.
@@
It is possible that for even thinner disks, perhaps an order of magnitude thinner in $h/r$, magnetic and thermal pressures remain permanently decoupled. Alternatively, disk dynamics may still depend on $h/r$, reaching a critical value where the disk can no longer retain its magnetic flux [55]. In that scenario, the disk would transition toward the NT state, theoretically contributing to BH spin-up. Observational evidence, such as jet quenching in the razor-thin XRB soft state, may support this idea. In contrast to thin MADs, which are dominated by super-thermal mainly turbulent magnetic fields at the midplane, accretion disks dominated by super-thermal toroidal fields (which is different from MADs) appear to be influenced by their thermal thickness, even if $p_{\mathrm{g}}$ is not dynamically relevant for the disk vertical equilibrium [89].
@@
We compute a simple fit for the dependence of $a_{\rm eq,MAD}$ on disk thickness, finding $a_{\rm eq,MAD}\simeq 0.31-2.7(h/r)^{2}.$ This provides a rough estimate of how $a_{\mathrm{eq}}$ might vary with the Eddington ratio, $\lambda_{\mathrm{Edd}}=\dot{m}/\dot{m}_{\mathrm{Edd}}$, and illustrates that the equilibrium spin plateaus at $a_{\mathrm{eq}}\lesssim 0.31$ in the limit of $h/r\to 0$. In future work, we plan to extend this model across a broader range of Eddington ratios.
VI.2 Comparison with other work
@@
Although our work is the first to measure BH spin-down for thin MADs with $h/r\leq 0.1$, previous studies have already examined jet efficiency in such disks [1, 84, 12]. For $h/r=0.1$, our results broadly agree with prior findings, showing similar jet efficiencies and magnetic fluxes at the event horizon. However, for even thinner disks ($h/r\sim 0.03$), our measured jet energy efficiencies differ significantly, with Scepi et al. [84] reporting jet efficiencies a factor of 5 lower. This discrepancy may stem from the differences in the measurement methods, as Scepi et al. [84] evaluate jet power at larger distances rather than the event horizon. Nonetheless, we note that our measurement of magnetic flux at the horizon, $\phi_{\mathrm{BH}}$, is consistent with Scepi et al. [84], even for MADs with $h/r\sim 0.03$.
@@
Recent work by Ricarte et al. [81] measured equilibrium spin in radiative MAD simulations but focused on thicker disks, with a minimum [^8] We note that our measurements of $h/r$ are not entirely comparable, as they use radiative and thermal pressure to measure $h/r$. $h/r\sim 0.18$. Direct comparison is difficult since our radiative simulations do not overlap in Eddington ratio, $\lambda_{\mathrm{Edd}}$, space. However, for $h/r\sim 0.18$, they report a higher equilibrium spin ($a_{\mathrm{eq}}\sim 0.8$) than what we find for $h/r=0.1$ ($a_{\mathrm{eq}}\sim 0.3$). Within our model, we expect $0.07\leq a_{\mathrm{eq}}\leq 0.3$ for such disks, in tension with their results.
@@
Ricarte et al. [81] also propose a general spin-down model applicable to all Eddington ratios, which we can compare to our radiative simulations at $\lambda_{\mathrm{Edd}}=0.35$ which have $h/r=0.03$ [51]. At this Eddington ratio they report an equilibrium spin of $a_{\mathrm{eq}}\sim 1$, whereas we find $a_{\mathrm{eq}}\sim 0.3$. Moreover, their results show no sign of convergence in $a_{\mathrm{eq}}$ for thinner disks; instead, their equilibrium spin continues increasing as disk thickness decreases, eventually reaching values consistent with the NT disk ($a_{\mathrm{eq}}=1$).
VI.3 Observational implications
@@
This work is timely for understanding BH spin measurements in x-ray binaries, where mass transfer–the process by which a donor star loses mass to a BH–is a key aspect of evolution and may also govern the BH spin evolution. Stellar-mass BH spins are directly measured via gravitational waves (GWs) in compact binary mergers [48, 101, 49, 15], constrained in XRBs using x-ray reflection [53, 80, 14] and continuum fitting [66, 90]. However, these methods yield conflicting spin distributions: GW observations show a distribution centered around $|a|\sim 0.2$, whereas EM observations find a distribution skewed toward higher BH spins, $a\gtrsim 0.5$. This discrepancy may arise from GW sources and XRBs belonging to distinct binary populations, which have different mass transfer histories [23, 20]. Spin evolution during mass transfer involves violent accretion episodes that can significantly alter the BH natal spin [77]. Whether these hypercritical accretion phases are linked to ultraluminous x-ray sources (ULXs) remains uncertain, but the mere fact that ULXs exist confirms that some binaries accrete far above Eddington [44]. The spin evolution during these highly super-Eddington accretion episodes has typically been modeled using the NT or Bardeen [2] theories, which predict BH spin evolution toward high spins–consistent with the BH spin distributions observed in high-mass XRBs [23, 77]. However, binaries can undergo diverse accretion spectral states, some of which launch powerful jets [13, 44]. For such jetted mass transfer phases, our MAD spin evolution model can provide a more accurate description. Moreover, because it predicts lower final spins, it may yield spin distributions that align better with those measured from GW sources–potentially helping to resolve the observed discrepancy.
@@
A leading model for long GRBs is the collapsar scenario, where a rapidly rotating massive star collapses into a BH, which launches powerful jets that produce the gamma-ray emission [98]. Our previous work on thick MADs [54] applied to collapsar GRBs demonstrated that their high accretion rates and strong jets make them ideal systems for BH spin-down [35, 99]. We showed that collapsar BHs can spin down to $a_{\mathrm{eq}}=0.07$ within seconds, well before the GRB ends. However, this work did not account for neutrino cooling [8], which alters the disk thickness and, consequently–as shown here–the equilibrium spin. Indeed, recent work by Issa et al. [32] finds that due to neutrino cooling of the accreting gas, collapsar BHs spin down to a higher equilibrium spin, $a_{\mathrm{eq}}\simeq 0.13$, and therefore likely fall between our cases of $h/r=0.1$ and $h/r=0.3$, to be consistent with our results. Their equilibrium spin measurement could serve as an initial condition for BH spin, which our model–better suited for binary evolution–could then evolve further in time.
@@
Jet feedback is crucial for galaxy formation and the evolution of large-scale cosmological structures like galaxy clusters. Understanding BH spin evolution in luminous quasars is essential, as the changes in BH spin directly impact the energy injection into the galaxy and cluster environment. Recent quasar formation simulations, which track accretion disk formation and evolution within a dynamically simulated galaxy, suggest the emergence of a super-thermal, toroidally dominated disk [30]. By interpolating the final snapshot of Hopkins et al. [30] simulation to H-AMR, Kaaz et al. [40] found that the central BH can accrete at super-Eddington rates, $\lambda_{\mathrm{Edd}}\equiv\dot{m}/\dot{m}_{\mathrm{Edd}}\sim 5$. Their disk settles into a MAD-like structure similar to the thin MADs studied here. Their simplified spin-down estimates suggest spin evolution on timescales comparable to the AGN duty cycle, $\sim 10\,\,\rm{Myr}$. This indicates that our results on BH spin-down are highly relevant for cosmological quasar accretion, as these systems accrete rapidly and exhibit MAD-like structures. This is especially important if one wants to also understand the implications of AGN BH spin measurements for the BH accretion and merger history [80].
@@
Observations of jetted accretion disk sources (XRBs and AGN) reveal a strong correlation between the jet luminosity ($L_{\mathrm{jet}}$) and accretion luminosity ($\dot{m}$), independent of accretor mass, and spanning a wide range of Eddington ratios ($\lambda_{\mathrm{Edd}}$) [85, 11, 59, 10, 103]. This suggests little variation in jet efficiency, $\eta_{\mathrm{EM}}=L_{\mathrm{jet}}/\dot{m}c^{2}$, across these systems. Our results show that in MADs, $\eta_{\mathrm{EM}}$ remains independent of disk thickness for $h/r\leq 0.1$, and even when comparing thick ($h/r=0.3$) and thin ($h/r=0.1$) disks, jet efficiency decreases by only a factor of 2 for fixed $a$. Such an astrophysically small variation, within a factor of $2$, in $\eta_{\mathrm{EM}}$ for fixed BH spin across different disk properties would be hard to distinguish observationally. This might explain the universal correlation between jet luminosity and accretion rate observed across diverse astrophysical systems over a wide range of Eddington ratios.
@@
Our results suggest that if all jetted BHs are in the MAD state, they should spin down to low spin values, $a_{\mathrm{eq}}\lesssim 0.3$, regardless of their Eddington ratio, given sufficient time. This has significant implications for the cosmological evolution of BHs in AGN, directly impacting cosmological BH feedback. Additionally, in binary systems, these findings are crucial for LISA and LVK science, offering a potential explanation for the observed BH spin distribution.
Acknowledgements.